OpenStudy (shylovesall):

How do you graph these two together? Y=-x+2 Y=3x-1

Nnesha (nnesha):

just like a simple linear equation graph first equation and then the 2nd one

Nnesha (nnesha):

y=mx+b slope intercept form equation where m is slope and b is y-intercept so start with y-intercept lets say the equation is y=4x+9 so i draw a point at 9 and then slope which is rise /run

Nnesha (nnesha):

the slope is 4 in that example which is same as \[\rm \frac{ 4 }{ 1 }=\frac{rise}{run}\] 4 =rise 1= run so go up 4 and then 1 to the right  and then draw a line passes through those two points